OpenStudy (anonymous):

Metal to who ever can help me catch up on my geometry work ive been in the hospital and need help please!!!!!??? FLVS

OpenStudy (anonymous):

Q?

OpenStudy (anonymous):

theres a couple

OpenStudy (anonymous):

yes....if i know it, will help

OpenStudy (anonymous):

idk if i got the first few right but i dont understand it at all.

OpenStudy (anonymous):

Are you still gonna help me @Isag

OpenStudy (anonymous):

iam not good at this, but let me try....

OpenStudy (anonymous):

alright thank you

OpenStudy (anonymous):

@***[ISURU]*** can you pls have a look at this problem

OpenStudy (whpalmer4):

1 is a parallelogram, which is more narrowly defined than trapezoid. Both pairs of sides are parallel in this figure. The angles are not all equal, however, so it is neither a square nor a rectangle, which are even more narrowly defined than a parallelogram.

OpenStudy (whpalmer4):

All squares are rectangles. All rectangles are parallelograms. All parallelograms are trapezoids.

OpenStudy (whpalmer4):

An equilateral triangle has 3 equal sides, and 3 equal angles. If you have 4 congruent sides, you are either a square or a rhombus. If the angles are also congruent, you are a square. I would use the distance formula to find the distance between two points, such as the endpoints of the sides on a quadrilateral. If they turn out to be equal, you have established that they are congruent. Find the slope of a line by taking the difference in the y values of the two points and dividing it by the difference in the x values of the two points (must be in the same order!). \[m = \frac{5-3}{4-(-2)} = \] For gross classification, slope of 0 = horizontal line. Undefined slope = vertical line. Positive slope = line that goes up and to the right. Negative slope = down and to the right. Parallel lines have equal slopes. Perpendicular lines have slopes whose product is -1.

OpenStudy (whpalmer4):

To find the point along a line that is ratio a:b from one end, calculate (a/(a+b)) and multiply that by the x and y values of the point on the left. Add the result to the corresponding x or y value to get the x or y value of the dividing point. The perimeter is just the sum of the lengths of the sides of the shape. Use the distance formula if needed to find their lengths. Distance formula for distance between two points \((x_1,y_1)\) and \((x_2,y_2)\) is\[d = \sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\] Formula for the area of a triangle is \[A = \frac{1}{2}b*h\]where \(h\) is height or altitude and \(b\) is length of base.

OpenStudy (whpalmer4):

I think that's all the background needed to knock off those problems. Good luck, especially with your recovery!

OpenStudy (anonymous):

got a 40% but thanks though it still helped. see even with help i still suck at geometry